OpenStudy (anonymous):

find all zeros. f(x)=x(5x-2)(x^2+1) PLEASE HELP ME STEP BY STEP.

OpenStudy (freckles):

you can find the zeros of f by setting f(x)=0 and solving for x you already have a factored expression for f Recall if a*b=0 then a=0 or b=0 or both=0 so just solve your three equations you have there: x=0 or 5x-2=0 or x^2+1=0 on of these equations is already solved

OpenStudy (anonymous):

ok so i do this 5x-2=0 / / 5 5

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

what does that mean?

OpenStudy (freckles):

are you dividing both sides by 5?

OpenStudy (freckles):

\[\frac{5x-2}{5}=\frac{0}{5} \\ \frac{5x}{5}-\frac{2}{5}=0 \\ \frac{\cancel{5}x}{\cancel{5}}-\frac{2}{5}=0 \\ x-\frac{2}{5}=0\] if so that is fine what next step would you perform

OpenStudy (anonymous):

2/5 = 0 0 0 x=2/5

OpenStudy (freckles):

right

OpenStudy (freckles):

well x=2/5 I think you meant you are just going to add 2/5 on both sides but yeah x=2/5 is right

OpenStudy (freckles):

now if you are to find the complex zeros you also need to solve x^2+1=0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

could you start me off

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

subtract one on both sides

OpenStudy (freckles):

x^2=-1 now take square root of both sides don't forget your plus or minus part

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

0^2?

OpenStudy (freckles):

the square root of -1 isn't 0 it is imaginary people normally call sqrt(-1) , i

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i and-i

OpenStudy (freckles):

yes

OpenStudy (anonymous):

so -1=-i (1)=i

OpenStudy (freckles):

x^2+1=0 (or the equivalent equation x^2=-1) gives the solutions x=-i or x=i since (-i)^2=(-1)^2(i)^2=1(-1)=-1 and i^2=-1

OpenStudy (anonymous):

ok thanks more?

OpenStudy (freckles):

more what?

OpenStudy (anonymous):

more help please

OpenStudy (freckles):

you can post a question if you want

OpenStudy (anonymous):

help me freckles